OpenStudy (anonymous):

Can i get help with this question on derivatives please please please with the cherry on top!...................

OpenStudy (johnweldon1993):

Haha sure :) what do you need help with?

OpenStudy (akashdeepdeb):

Ask away!

OpenStudy (anonymous):

thanks @johnweldon1993 and @AkashdeepDeb , i'm stuck on this question, i can't go any further:(

myininaya (myininaya):

can barely read it but I think I made out the function \[f(x)=\frac{\arctan(x)}{x^4-1} \]

OpenStudy (akashdeepdeb):

\[f(x) = \frac{\tan^{-1} x}{x^4 - 1} ~~?\]

OpenStudy (johnweldon1993):

^That's what I'm seeing too, can you confirm @study17 ?

myininaya (myininaya):

It looks like you know the quotient rule but you didn't write the derivative of the bottom next to the arctan(x) part .

OpenStudy (anonymous):

lol sorry for my bad writing @johnweldon1993 yes @AkashdeepDeb is right, that's what i mean

myininaya (myininaya):

\[f'(x)=\frac{[\arctan(x)]' (x^4-1)-\arctan(x) [(x^4-1)]'}{(x^4-1)^2}\]

OpenStudy (johnweldon1993):

Right, as myininaya has stated, when you did the quotient rule, you did not take the derivative of the bottom \[\large f'(x) = \frac{u'v - uv'}{v^2}\] That v' up there is where you just wrote in 'v' without taking its derivative

OpenStudy (johnweldon1993):

Also notice in the denominator, it is \(\large v^2\) not \(\large v' ^2\)

OpenStudy (anonymous):

oh crap yeah thanks everyone i see where i went wrong now, i was doing this in the car lol... this is what i got when i redid it but i honestly have no idea what to do now

OpenStudy (akashdeepdeb):

To simplify things too, write \((x^4 - 1) = (x^2-1)(x^2+1)\) and then simplify.

OpenStudy (anonymous):

this is the answer on the back of my book, i just really don't know how to get there, do you mind showing me step by step please if you don't mind

OpenStudy (johnweldon1993):

Ahh perfect I was simplifying it and that's exactly what I got lol...hang on 1 sec

OpenStudy (johnweldon1993):

\[\large f'(x) = \frac{\frac{x^4 - 1}{1 + x^2} - 4x^3\tan^{-1}(x)}{(x^4 - 1)^2}\] Lets get a common denominator in the top part, which will be \(\large 1 + x^2\) So we need to multiply the \(\large -4x^3tan^{-1}(x)\) by that Which gives us \[\large \frac{\frac{x^4 - 1 - ((4x^3 \tan^{-1}(x) + (4x^5 \tan^{-1}(x))}{1 + x^2}}{(x^4 - 1)^2}\] Now, if we turn this to multiplication, fip the bottom fraction and multiply \[\large \frac{x^4 - 1 - 4x^3 \tan^{-1}(x) -4x^5 \tan^{-1}(x))}{(1 + x^2)(x^4 - 1)^2}\] Now, if we factor out a \(\large -4x^3 \tan^{-1}(x)\) we would have \[\large \frac{(x^4 - 1) - (4x^3 \tan^{-1}(x))(1 + x^2)}{(1 + x^2)(x^4 - 1)^2}\] Break this into 2 fractions \[\large \frac{x^4 - 1}{(1 + x^2)(x^4 - 1)^2} - \frac{4x^3 \tan^{-1}(x)(1 + x^2)}{(1 + x^2)(x^4 - 1)^2}\] Notice anything? Like \[\large \frac{\cancel{x^4 - 1}}{(1 + x^2)(x^4 - 1)\cancel{^2}} - \frac{4x^3 \tan^{-1}(x)\cancel{(1 + x^2})}{\cancel{(1 + x^2)}(x^4 - 1)^2}\] So all we have left is \[\large \frac{1}{(1 + x^2)(x^4 - 1)} - \frac{4x^3 \tan^{-1}(x)}{(x^4 - 1)^2}\]

OpenStudy (johnweldon1993):

Sorry that took so long >.< lol

OpenStudy (anonymous):

wow thanks you've been such a big help!! i'll look through this now and see if i get it completely and see if i have any problems understanding it, thanks so much!

OpenStudy (anonymous):

@johnweldon1993 i'm stuck on simplifying again lol, help please genius!

OpenStudy (johnweldon1993):

Alright so right after you figure out all the derivatives and plug them in..what you should have is \[\large \frac{(x^2 + 1)(-\frac{1}{\sqrt{9 - x^2}}) - 2xcos^{-1}(\frac{x}{3})}{(x^2+1)^2}\] which you do so now okay lets multiply that first pat like you did...ut 1 thing is that when you multiply that...you get \[\large \frac{(\frac{-x^2 - 1}{\sqrt{9 - x^2}}) - 2xcos^{-1}(\frac{x}{3})}{(x^2+1)^2}\] You didnt multiply through the negative sign...alright so let me look for a little bit more :)

OpenStudy (johnweldon1993):

Lets see, does the answer in the back of the book look like \[\large -\frac{1}{(x^2 + 1)(\sqrt{9 - x^2}} - \frac{2x \cos^{-1}(\frac{x}{3})}{(x^2 + 1)^2}\] or do I have to keep working on it? lol

OpenStudy (anonymous):

That is exactly correct except in the back of the book there's no - sign in front of the answer but i think the back of the book is wrong (it's been wrong a lot of times before) so you're right yeah :)

OpenStudy (johnweldon1993):

Hmm, I'll have to look it over, but I'll do it as I post here for you :)

OpenStudy (johnweldon1993):

\[\large \frac{\frac{-(x^2 + 1)}{\sqrt{9 - x^2}} - 2x\cos^{-1}(\frac{x}{3})}{(x^2 + 1)^2 }\] So same concept here...we need the common denominator in the top part....which will be the \(\large \sqrt{9 - x^2}\) and this tme I'll just write it like you pointed out last time :P \[\large \frac{\frac{-(x^2 + 1)- 2x\cos^{-1}(\frac{x}{3})\sqrt{9 - x^2}}{\sqrt{9 - x^2}}}{(x^2 + 1)^2 }\] So now flip the bottom fraction...and turn this to multiplication \[\large \frac{-(x^2 + 1) - 2x\cos^{-1}(\frac{x}{3})\sqrt{9 - x^2}}{(x^2 + 1)^2(\sqrt{9 - x^2})}\] So lets turn this again into 2 fractions \[\large \frac{-(x^2 + 1)}{(x^2 + 1)^2(\sqrt{9 - x^2})} - \frac{2x\cos^{-1}(\frac{x}{3})\sqrt{9 - x^2}}{(x^2 + 1)^2(\sqrt{9 - x^2})}\] And now we can see what stuff cancels \[\large \frac{-\cancel{(x^2 + 1)}}{(x^2 + 1)\cancel{^2}(\sqrt{9 - x^2})} - \frac{2x\cos^{-1}(\frac{x}{3})\cancel{\sqrt{9 - x^2}}}{(x^2 + 1)^2(\cancel{\sqrt{9 - x^2})}}\] Which leaves us with \[\large -\frac{1}{(x^2 + 1)(\sqrt{9 - x^2})} - \frac{2x\cos^{-1}(\frac{x}{3})}{(x^2 + 1)^2}\] Yeah still have that negative sign there :)

OpenStudy (anonymous):

you know on the first line? why is the common denominator only the square root of 9-x^2, i thought 3 would be a common denominator as well because we have x/3 ?

OpenStudy (johnweldon1993):

Well remember \[\large \cos^{-1}(\frac{x}{3}) \cancel{=} \frac{\cos^{-1}(x)}{3}\]

OpenStudy (anonymous):

oh i see i see! on the 2nd line i don't understand how the square root of 9-x^2, ended up as a denominator as well, is it not meant to cancel out like the way our denominator of 1+x^2 cancelled out in the 1st question i asked?

OpenStudy (johnweldon1993):

It didn't cancel out right away in the first question :) So the first line was \[\large \frac{\frac{-(x^2 + 1)}{\sqrt{9 - x^2}} - 2x\cos^{-1}(\frac{x}{3})}{(x^2 + 1)^2 }\] And lets just focus on the top part because that is what your question is regarding...so \[\large \frac{-(x^2 + 1)}{\sqrt{9 - x^2}} - 2x\cos^{-1}(\frac{x}{3})\] in order to subtract those 2 things....we need a common denominator...which we know now is \(\large \sqrt{9 - x^2}\) right? so notice that the \(\large 2x\cos^{-1}(\frac{x}{3})\) needs to have that as its denominator...so we need to multiply it by \(\large \frac{\sqrt{(9 - x^2)}}{\sqrt{9 - x^2)}}\) right? so now we'll have \[\large \frac{-(x^2 + 1)}{\sqrt{9 - x^2}} - 2x\cos^{-1}(\frac{x}{3}) \times \frac{\sqrt{(9 - x^2)}}{\sqrt{9 - x^2)}} \] *Notice it is ONLY being multiplied to the \(\large 2x\cos^{-1} (\frac{x}{3})\) Now that will become \[\large \frac{-(x^2 + 1)}{\sqrt{9 - x^2}} - \frac{2x\cos^{-1}(\frac{x}{3})\sqrt{9 - x^2}}{\sqrt{9 - x^2 }} \] Right? and because its redundant to write it like that...we combine the 2 over the common denominator \[\large \frac{-(x^2 + 1) - 2x\cos^{-1}(\frac{x}{3})\sqrt{9 - x^2}}{\sqrt{9 - x^2}}\]

OpenStudy (anonymous):

yeah i get that part, i just don't know how the square root of 9-x^2 ended up being beside (x^2+1)^2 ??

OpenStudy (johnweldon1993):

Okay that was the next part...so lets bring back the whole thing...the 2nd line \[\large \frac{\frac{-(x^2 + 1)- 2x\cos^{-1}(\frac{x}{3})\sqrt{9 - x^2}}{\sqrt{9 - x^2}}}{(x^2 + 1)^2 }\] So this is where we flip the bottom fraction....and no...the bottom fraction is not what you may think at first glance...think of this as \[\large \frac{\frac{-(x^2 + 1)- 2x\cos^{-1}(\frac{x}{3})\sqrt{9 - x^2}}{\sqrt{9 - x^2}}}{(x^2 + 1)^2 }\] \[\large \frac{\frac{-(x^2 + 1) - 2x\cos^{-1}(\frac{x}{3})\sqrt{9 - x^2}}{\sqrt{9 - x^2}}}{\frac{(x^2 + 1)^2}{1}}\] so when I say here to fip the bottom fraction...and muliply...what I mean is Flip \(\large \frac{(x^2 + 1)^2}{1} \) to become \(\large \frac{1}{(x^2 + 1)^2}\) and then multiply this to the top part of the whole fraction...so now we have \[\large \frac{-(x^2 + 1)^2 - 2x\cos^{-1}(\frac{x}{3})\sqrt{9 - x^2}}{(x^2 + 1)^2\sqrt{9 - x^2}}\]

OpenStudy (anonymous):

thanks so much @johnweldon1993 you're amazing!! i understand it completely!! i have another problem though..yet again lol

OpenStudy (johnweldon1993):

Still having trouble carrying over the common denominator :) you did it perfect until right \[\large \frac{\frac{x^2}{\sqrt{1 - x^2}} - 2x\sin^{-1}(x)}{(x^2)^2}\] So we'll focus on the top part, like you did And you did get \[\large \frac{x^2 - 2x\sin^{-1}(x)\sqrt{1 - x^2}}{(\sqrt{1 - x^2})}\] But remember..that was ust focusing on the TOP part...we ignored the \(\large (x^2)^2\) part....so what we have here is \[\large \frac{x^2 - 2x\sin^{-1}(x)\sqrt{1 - x^2}}{(\sqrt{1 - x^2})}\] \[\large \frac{\frac{x^2 - 2x\sin^{-1}(x)\sqrt{1 - x^2}}{\sqrt{1 - x^2}}}{\frac{(x^2)^2}{1}}\] remember that part? now we flip and multiply

OpenStudy (anonymous):

oh yeah! ok so i applied that and i'm going the right way but i'm stll not completely right;(

OpenStudy (johnweldon1993):

From what I see EVERYTHING is correct...except maybe 1 thing is where you have \[\large \frac{x^{-2}}{\sqrt{1 - x^2}}\] This will be why you see I didnt write \(\large x^4\) but instead always wrote \(\large (x^2)^2\) So what you should have is \[\large \frac{x^2}{(x^2)^2 \sqrt{1 - x^2}} - \frac{2x\sin^{-1}(x) }{(x^2)^2}\] and notice \[\large \frac{\cancel{x^2}}{(x^2)\cancel{^2} \sqrt{1 - x^2}} - \frac{2x\sin^{-1}(x) }{(x^2)^2}\] Which would make your final answer \[\large \frac{1}{(x^2) \sqrt{1 - x^2}} - \frac{2x\sin^{-1}(x) }{(x^2)^2}\] does that look better?

OpenStudy (anonymous):

oh i see i see, yeas thats like the back of the book except that in the back of my book there's no x beside 2 in 2xsin^-1(x) and it's denominater is x^3 instead of (x^2)^2 but then again the back of the book has made lots of mistakes before:)

OpenStudy (johnweldon1993):

Well no we can make it that... we can write \(\large (x^2)^2\) as \(\large x^4\) and then we'll have \[\large \frac{1}{(x^2) \sqrt{1 - x^2}} - \frac{2x\sin^{-1}(x) }{(x^4)}\] Now notice if we do that, we can cross out the 1 x in the top part...and cross out 1 'x' in the bottom...leaving us with \[\large \frac{1}{(x^2) \sqrt{1 - x^2}} - \frac{2\sin^{-1}(x) }{(x^3)}\]

OpenStudy (anonymous):

oh yeah very clever! now i'm stuck on a new type of question lmao

OpenStudy (johnweldon1993):

Nope nope nope....lol :P we ignore the 1/4 until the end \[\large \frac{d}{dx} sin^{-1} (x^3)\] That's what this is right? the x is raised to the 3rd power?

OpenStudy (anonymous):

oh!! yep the x is raise to the 3rd power :)

OpenStudy (johnweldon1993):

So now...what is the derivative of that?

OpenStudy (anonymous):

is it\[\frac{ 3x^{2} }{ \sqrt{1-x^{5}} }\]

OpenStudy (johnweldon1993):

Soooo close derivative of \(\large sin^{-1}(x) = \frac{1}{\sqrt{1 - x^2}}\) Sooo if our 'x' is actually x^3....we have \(\large sin^{-1}(x) = \frac{1}{\sqrt{1 - (x^3)^2}}\) And remember that when we multiply exponents through parenthesis...we multiply...not add...so we have \(\large \frac{3x^2}{\sqrt{1 - x^6}}\)

OpenStudy (anonymous):

omg lol i make the stupidest mistakes!! ok so i subbed in the 1/4 now and i got \[\frac{ \frac{ 3 }{ 16} }{\frac{ 3\sqrt{455}}{ 64} }\] and i don't know what to do from there..

OpenStudy (anonymous):

do i flip the bottome denominator and multiply 3 by 64 and 16 by 3 root of 455?

OpenStudy (johnweldon1993):

Indeed^ :)

OpenStudy (anonymous):

at the back of the book it's 4 sqare root 455 over 455 though?

OpenStudy (johnweldon1993):

So lets see \[\large \frac{64 \times 3}{16 \times 3\sqrt{455}} = \frac{192}{48\sqrt{455}} = \frac{4}{\sqrt{455}}\] Are you okay up to here? just 1 more step...we just need to rationalize the denominator

OpenStudy (anonymous):

i forgot how to rationalize denominators especially at 2 am lol

OpenStudy (johnweldon1993):

Oh you poor thing!! lol You just multiply this like \[\large \frac{4}{\sqrt{455}} \times \frac{\sqrt{455}}{\sqrt{455}}\] which then equals \[\large \frac{4\sqrt{455}}{455} \]

OpenStudy (anonymous):

i know right lol!! oh yeah!! hey i'm doing this right, right?

OpenStudy (johnweldon1993):

Perfect :)

OpenStudy (anonymous):

i don't know if i got g(x) right and i'm a little confused on it:(

OpenStudy (johnweldon1993):

So this one is just finding the derivative right?

OpenStudy (anonymous):

yep:)

OpenStudy (johnweldon1993):

Okay so lets do it out The derivative as you wrote of \(\large \tan^{-1}(x) = \frac{1}{1 + x^2}\) And so like the others we would have \[\large \frac{1}{1 + (\frac{x}{2})^2} \times \frac{1}{2}\] Notice that it is times 1/2 because the derivative of the inside (the x/2) would be 1/2) so \[\large \frac{1}{2(1 + (\frac{x^2}{4}))}\]

OpenStudy (anonymous):

okay thanks so much john you've been the best help ever! good night!

OpenStudy (johnweldon1993):

Yeah get some sleep hun :) If you need anything else just let me now :)

OpenStudy (johnweldon1993):

Know* terrible grammar XD